OpenStudy (anonymous):

What is the slope of the line that contains the points (-3, -2) and (4, -2)?

OpenStudy (anonymous):

\[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] m is slope

OpenStudy (anonymous):

\[(-3, -2)\implies (x_1,y_1)\] \[(4, -2) \implies (x_2,y_2)\]

OpenStudy (anonymous):

- 4 undefined - 4/7 0

OpenStudy (anonymous):

K I gave you all the information you need now do it yourself.

OpenStudy (anonymous):

well duh im trying to learn it...........so its -4/7

OpenStudy (anonymous):

Show me how you got that

OpenStudy (anonymous):

Seems like you just took a guess

OpenStudy (anonymous):

taking a guess is not my style hero........ -2-2/4+3.......just trying man so -4/7

OpenStudy (anonymous):

Naw, wrong again, I'll set it up for ya \[m=\frac{ -2-(-2) }{ 4-(-3) } =?\]

OpenStudy (anonymous):

You see now?

OpenStudy (anonymous):

4/7 but thats not one of the answers

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

its zero

OpenStudy (anonymous):

How are you getting 4/7 I just set it up for you, ok lets look the numerator, -2-(-2), drop the brackets negative * negative = positive -2+2 = 0 :P

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

wait so is it just 0 or 0/7

OpenStudy (anonymous):

0/7 = 0, 0 divided by anything always equals 0, and when 0 is in the denominator you get undefined.

OpenStudy (anonymous):

so is the answer 0 or undefined they are both listed

OpenStudy (anonymous):

lol it's 0

OpenStudy (anonymous):

\[\frac{ 0 }{ x } = 0~~~~ \frac{ x }{ 0 } = undefined \]

OpenStudy (anonymous):

ok sorry I'm just am stupid at math .........i have pretty much strait A's and then there is math

OpenStudy (anonymous):

What is the slope of the line that contains the points (2, 6) and (-1, -6)? so 4